OpenStudy (anonymous):

inverse of: f(x) = ln(2x)-3 :S

OpenStudy (anonymous):

well here again, switch the x and y's think of f(x) = y b.t.w. x=ln(2y)-3 x+3=ln(2y) e^(ln2y)=e^(x+3) 2y=e^(x+3) y=(1/2)e^(x+3) .. and that is your inverse

OpenStudy (anonymous):

I get through till your third step.

OpenStudy (anonymous):

see that again, is just a rule.. it will be in your text book so remember this.. forever! e^(lnx) = x

OpenStudy (anonymous):

because you can't just divide by an ln. it doesn't work like that